If a cylinder has a raius of 3 cm and a length of 10 cm. and every dimension of the cylinder is multiplied by 3 to form a new cylinder. How is the ratio of the volumes related to the ratio of the corresponding dimensions?
Answer by Alan3354(69443)

If a cylinder has a raius of 3 cm and a length of 10 cm. and every dimension of the cylinder is multiplied by 3 to form a new cylinder. How is the ratio of the volumes related to the ratio of the corresponding dimensions?
--------------------
The volume is directly related to the square of the radius, and linearly related to the length.
V1 = pi*r^2*h
V2 = 27*V1
---------
V = k*r^2 directly related to the square of the radius
V = kh directly related to the length